# 0.5 eV QCD Axion Cosmology

### v47

@@ Abstract
A simple yet compelling physical picture is proposed for the nature of dark matter, dark energy, and the Big Bang. The proposal leads to a prediction with 0.53 parts per million precision for the 0.5 eV rest-mass energy of the quantum chromodynamic (QCD) axions that form the dark matter within this picture. A straightforward resolution of the Hubble tension within this picture follows from the agreement between the prediction for the present expansion rate of the universe and the best present determination from observations of the late universe.
@@
show]nbrayali@google.com
@@ 1 Introduction
$H_{0}=71.931~(8)~{\rm km\cdot sec^{-1}\cdot Mpc^{-1}}~[{\rm 0.11~per~mille}].$ (1)
@@
This combines with the percent-level precise prediction for the strength of the axion-photon coupling (See Appendix D),
@@
In this historical context, the key insight is simply to regard the particle-antiparticle pairs, which are responsible for the dark energy described by the cosmological constant, as forming the dark matter in the universe. The Big Bang, then, is just the moment when massless matter and antimatter bind to form the dark matter, in the form of these particle-antiparticle pairs. As the universe expands, the number density of dark matter, $n_{A}(t)\propto 1/a(t)^{3}$, dilutes with the volume of the expanding universe given by the scale factor of the universe, $a(t)$, at time $t$ after the Bang, and, thus, the kinetic energy of the massless matter and antimatter inside the dark matter, $E(t)\propto a(t)$ must grow to keep constant in time the value of the cosmological constant, $\Lambda\propto n_{A}(t)E(t)^{3}$.
@@
where the best present estimate for the topological susceptibility of the QCD vacuum from Chiral Perturbation Theory (CPT) is $\sqrt{\chi_{\rm QCD}}=5690~(50)~{\rm MeV^{2}}~[{\rm 8.8~per~mille}]$ (M. Gorghetto & G. Villadoro, 2019).
@@
where the exact value of the astronomical unit ${\rm AU}=149~597~870.700~{\rm km}$ (N. Capitaine et al., 2012) was used to derive the numerically exact value of the cosmological time, $h/H_{0}=({\rm AU}/100)(180/\pi)(60)^{2}\times 10^{6}~{\rm sec}=3.085~677~581\times 10^{17}~{\rm sec}.$
@@
$\Omega_{\Lambda}({\rm CPT})$ $=$ $\left(0.35\right)\left(\frac{107.6~(1.6)}{104.2}\right)^{6}\frac{1}{(0.714~405)^{2}}$ (20)
$=$ $0.831~(75)~[90~{\rm per~mille}].$
@@
To push the precision of the first principles estimate of the cosmological constant and to determine the last remaining cosmological parameter, $\Omega_{B}h^{2}$, that of the baryons, we must leverage the numerically exact value for the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}=0.399~286~(14)$, given in Eq. (5), and derived in Sec. 4:
@@
where the value of $\sqrt{\chi_{\rm QCD}}$ agrees with the best present estimate from chiral perturbation theory (CPT) $\sqrt{\chi_{\rm QCD}}({\rm CPT})=5~690~(50)~{\rm MeV}^{2}~[8.8~{\rm per~mille}]$ (M. Gorghetto & G. Villadoro, 2019), to within two percent, but is nearly 26000 times more precise. Repeating the quick and dirty calculation of the cosmological constant in Eq. (20), but now with the Big Bang kinetic energy, $\check{E}=x_{k}kT_{\pi}=104.143~568~(18)~[0.17~{\rm ppm}]$, given by the first principles prediction for the Big Bang temperature in Eq. (3), yields the value of the cosmological parameter for the cosmological constant, $\Omega_{\Lambda}h^{2}$, to 22 ppm (See Appendix B for the derivation of the precision of the estimate):
$\Omega_{\Lambda}h^{2}$ $=$ $\left(0.35\right)\left(\frac{104.143~568}{104.218~7}\right)^{6}$ (27)
$=$ $0.352~362~8~(78)~[{\rm 22~ppm}].$
@@ 9 Conclusions
Among astrophysicists during the past forty years, three main “pillars” have come to be viewed as supporting the now-standard cold dark matter cosmology ($\rm\Lambda CDM$}): the anisotropy of the cosmic microwave background (CMB), the “primordial” abundance of chemical elements with low atomic mass whose nuclei were synthesized in the Big Bang (BBN), and the redshift-magnitude relation of Type Ia supernovae (P. J. E. Peebles, 2020, pg. 323–340). The physical picture of dark matter, dark energy, and the Big Bang, presented in this Letter, rests on the same “pillars” of support, but does so more naturally. In particular, the proposed physical picture naturally resolves the $5.0\sigma$ tension in the present expansion rate of the universe between $\rm\Lambda CDM$} and “late” universe observations, and does so in favor of the observations.
@@
Comparing the last line of Eq. (A2) with the expression for the graviational self-energy in Eq. (8), we find the value of the dimensionless constant, $N_{U}=1/(2\pi^{2})$. We emphasize that this result requires that the 0.5 eV QCD axion dark matter be a spin-less bound-state of particle and anti-particle, and, thus, the axion must obey Bose-Einstein statistics (L. D. Landau & E. M. Lifshitz, 1958, pg. 205), just as photons do (P. J. E. Peebles, 1993, pg. 134).
@@
Appendix D 0.5 eV QCD Axion-Photon Coupling Strength
@@
$g_{A\gamma\gamma}$ $=$ $C_{A\gamma}\frac{\alpha}{2\pi}\frac{1}{f_{A}}$ (D1)
@@
$C_{A\gamma}$ $=$ $\frac{E}{N}-\frac{2}{3}\frac{4m_{d}+m_{u}}{m_{d}+m_{u}}$ (D2)
@@
where the value of the anomaly ratio is set by the fact that all twelve flavors of “normal” quarks and leptons within the standard model of particle physics transform with the same axial charge, $Q_{A}=+1$, under the continuous axial $U(1)_{A}$ phase symmetry, $U_{A}(\alpha)=\exp(iQ_{A}\alpha)$, while the opposite axial charge, $Q_{A}=-1$, holds for “mirror” quarks and leptons which have the same correlation between chirality and helicity as “normal” antimatter (Right-handed mirror matter feels the weak nuclear force but left-handed mirror matter does not feel it; See L. M. Lederman & D. N. Schramm (1989), pg. 121–123): The 0.5 eV QCD axion dark matter, $A_{M}$, with a given flavor, $M$, is the unique quantum state formed by superposing the bound state of mirror matter and normal antimatter with the bound state of normal matter and mirror antimatter, such that the particle and the antiparticle with flavor $M$ inside the bound states have the same helicity and chirality, and such that the resulting quantum state for the axion, and hence the axion field, $\phi_{A}$, has the same discrete space-time symmetry properties as the dot product, ${\bf E}\cdot{\bf B}$, ensuring that the Lagrangian density, $\mathcal{L}_{A\gamma}$, is invariant under these transformations (charge-conjugation, space-inversion, and time-reversal; See L. M. Lederman & D. N. Schramm (1989), pg. 121–123). Finally, we combine Eq. (D1) and Eq. (D2) to find the value for the axion-photon coupling strength given in Eq. (7).

### v48

@@ Abstract
A simple yet compelling physical picture is proposed for the nature of dark matter, dark energy, and the Big Bang. The proposal leads to a prediction with 0.53 parts per million precision for the 0.5 eV rest-mass energy of the quantum chromodynamic (QCD) axions that form the dark matter within this picture. A straightforward resolution of the Hubble tension follows within this picture from the agreement between the prediction for the present expansion rate of the universe and the best present determination from observations of the late universe.
@@
show]nbrayali@gmail.com
@@ 1 Introduction
$H_{0}=71.939~5~(76)~{\rm km\cdot sec^{-1}\cdot Mpc^{-1}}~[{\rm 0.11~per~mille}].$ (1)
@@
This combines with the percent-level precise prediction for the strength of the axion-photon coupling (See Appendix E),
@@
In this historical context, the key insight is simply to regard the particle-antiparticle pairs, which are responsible for the dark energy described by the cosmological constant, as forming the dark matter in the universe. The Big Bang, then, is just the moment when massless matter and antimatter bind to form the dark matter, in the form of these particle-antiparticle pairs. As the universe expands, the number density of dark matter, $n_{A}(t)\propto 1/a(t)^{3}$, dilutes with the volume of the expanding universe given by the scale factor of the universe, $a(t)$, at time $t$ after the Bang, and, thus, the kinetic energy of the massless matter and antimatter inside the dark matter, $E(t)\propto a(t)$, must grow to keep constant in time the value of the cosmological constant, $\Lambda\propto n_{A}(t)E(t)^{3}$.
@@
where the best present estimate for the topological susceptibility of the QCD vacuum from Chiral Perturbation Theory (CPT) is $\sqrt{\chi_{\rm QCD}}=5~690~(50)~{\rm MeV^{2}}~[{\rm 8.8~per~mille}]$ (M. Gorghetto & G. Villadoro, 2019).
@@
where the exact value of the astronomical unit ${\rm AU}=149~597~870.700~{\rm km}$ (N. Capitaine et al., 2012) was used to derive the numerically exact value of the cosmological time, $h/H_{0}=({\rm AU}/100~{\rm km})(180/\pi)(60)^{2}\times 10^{6}~{\rm sec}=3.085~677~581\times 10^{17}~{\rm sec}.$
@@
$\Omega_{\Lambda}({\rm CPT})$ $=$ $\left(0.35\right)\left(\frac{107.6~(1.6)}{104.2}\right)^{6}\frac{1}{(0.719~395)^{2}}$ (20)
$=$ $0.820~(73)~[90~{\rm per~mille}].$
@@
To push the precision of the first principles estimate of the cosmological constant and to determine the last remaining cosmological parameter, $\Omega_{B}h^{2}$, that of the baryons, we must leverage the numerically exact value for the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}=0.399~286~(14)~[35~{\rm ppm}]$, given in Eq. (5), and derived in Sec. 4:
@@
where the value of $\sqrt{\chi_{\rm QCD}}$ agrees with the best present estimate from chiral perturbation theory (CPT) $\sqrt{\chi_{\rm QCD}}({\rm CPT})=5~690~(50)~{\rm MeV}^{2}~[8.8~{\rm per~mille}]$ (M. Gorghetto & G. Villadoro, 2019), to within two percent, but is nearly 26000 times more precise. At last, combining the Big Bang temperature from Eq. (3) with the 0.5 eV QCD axion rest-mass energy from Eq. (6) yields the cosmological parameter for the cosmological constant, $\Omega_{\Lambda}h^{2}$, to 22 ppm (See Appendix B for the derivation of the precision of the estimate and Appendix D for the derivation of the estimate itself):
$\Omega_{\Lambda}h^{2}$ $=$ $\frac{16\zeta(3)x_{k}^{6}y_{T}^{10}}{\pi^{3}}\frac{\left(m_{A}c^{2}\right)^{5}}{kT_{\pi}E_{\rm Pl}^{2}}\left(\frac{h}{H_{0}}\frac{1}{\hbar}\right)^{2}$ (27)
$=$ $0.352~460~0~(78)~[{\rm 22~ppm}],$
@@ 9 Conclusions
Among astrophysicists during the past forty years, three main “pillars” have come to be viewed as supporting the now-standard cold dark matter cosmology ($\rm\Lambda CDM$}): the anisotropy of the cosmic microwave background (CMB), the “primordial” abundance of chemical elements with low atomic mass whose nuclei were synthesized in the Big Bang (BBN), and the redshift-magnitude relation of Type Ia supernovae (P. J. E. Peebles, 2020, pg. 323–340). The physical picture of dark matter, dark energy, and the Big Bang, presented in this Letter, rests on the same “pillars” of support, but does so more naturally. In particular, the proposed physical picture naturally resolves the now more than $5\sigma$ tension in the present expansion rate of the universe between $\rm\Lambda CDM$} and “late” universe observations, and does so in favor of the observations.
@@
Comparing the last line of Eq. (A2) with the expression for the gravitational self-energy in Eq. (8), we find the value of the dimensionless constant, $N_{U}=1/(2\pi^{2})$. We emphasize that this result requires that the 0.5 eV QCD axion dark matter be a spin-less bound-state of particle and anti-particle, and, thus, the axion must obey Bose-Einstein statistics (L. D. Landau & E. M. Lifshitz, 1958, pg. 205), just as photons do (P. J. E. Peebles, 1993, pg. 134).
@@
Appendix D 0.5 eV QCD Axion Dark Energy
The derivation of the expression for the cosmological parameter of the cosmological constant, $\Omega_{\Lambda}h^{2}$, in Eq. (27), begins with the physical picture for the nature of dark energy presented in Section 2, but evaluated now at future infinity instead of the Big Bang:
$\Omega_{\Lambda}h^{2}$ $=$ $\frac{8\pi G}{3c^{2}}\left(\frac{h}{H_{0}}\right)^{2}\hat{U}_{\Lambda}\hat{n}_{A}$ (D1)
$=$ $\frac{8\pi}{3}\frac{\hbar c^{3}}{E_{\rm Pl}^{2}}\left(\frac{h_{0}}{H}\right)^{2}\left(N_{U}f_{A}\right)\left(N_{n}\left(\frac{\check{E}}{\hbar c}\right)^{3}\left(\frac{\check{E}}{\hat{E}}\right)^{3}\right)$
$=$ $\frac{8\pi}{3}N_{U}N_{n}\left(\frac{\check{E}^{9}}{f_{A}^{5}E_{\rm Pl}^{2}}\right)\left(\frac{m_{A}c^{2}}{m_{A}c^{2}}\right)^{5}\left(\frac{h}{H_{0}}\frac{1}{\hbar}\right)^{2}$
$=$ $\frac{8\pi}{3}N_{U}N_{n}x_{k}^{9}y_{T}^{10}\frac{\left(m_{A}c^{2}\right)^{5}}{kT_{\pi}E_{\rm Pl}^{2}}\left(\frac{h}{H_{0}}\frac{1}{\hbar}\right)^{2}$
$=$ $\frac{16\zeta(3)x_{k}^{6}y_{T}^{10}}{\pi^{3}}\frac{\left(m_{A}c^{2}\right)^{5}}{kT_{\pi}E_{\rm Pl}^{2}}\left(\frac{h}{H_{0}}\frac{1}{\hbar}\right)^{2},$
where $\hat{U}_{\Lambda}=N_{U}\hat{E}^{3}/E_{\rm Pl}^{2}=N_{U}f_{A}$ is the dark energy inside each axion at future infinity, and $\hat{n}_{A}=(\check{\ell}/\hat{\ell})^{3}(T_{\pi}/T_{\gamma})^{3}n_{A}=N_{n}(\check{E}/\hbar c)^{3}(\check{E}/\hat{E})^{3}$ is the axion number density at future infinity. See Appendix A for the derivation of the dimensionless constants, $N_{U}=1/(2\pi^{2})$ and $N_{n}=12\zeta(3)/(\pi^{2}x_{k}^{3})$, but see Appendix C for the conformal scale factors, $\hat{\ell}=\hat{E}/f_{A}$ and $\check{\ell}=\check{E}/f_{A}$. The see-saw relation, $\check{E}\hat{E}=f_{A}^{2}$ (See Section 2), was used in the third line, while the relations, $\check{E}=x_{k}kT_{\pi}$, and, $m_{A}c^{2}f_{A}=\sqrt{\chi_{\rm QCD}}=(kT_{\pi}/y_{T})^{2}$, were used in the fourth line.
Appendix E 0.5 eV QCD Axion-Photon Coupling Strength
@@
$g_{A\gamma\gamma}$ $=$ $C_{A\gamma}\frac{\alpha}{2\pi}\frac{1}{f_{A}}$ (E1)
@@
$C_{A\gamma}$ $=$ $\frac{E}{N}-\frac{2}{3}\frac{4m_{d}+m_{u}}{m_{d}+m_{u}}$ (E2)
@@
where the value of the anomaly ratio is set by the fact that all twelve flavors of “normal” quarks and leptons within the standard model of particle physics transform with the same axial charge, $Q_{A}=+1$, under the continuous axial $U(1)_{A}$ phase symmetry, $U_{A}(\alpha)=\exp(iQ_{A}\alpha)$, while the opposite axial charge, $Q_{A}=-1$, holds for “mirror” quarks and leptons which have the same correlation between chirality and helicity as “normal” antimatter (Right-handed mirror matter feels the weak nuclear force but left-handed mirror matter does not feel it; See L. M. Lederman & D. N. Schramm (1989), pg. 121–123): The 0.5 eV QCD axion dark matter, $A_{M}$, with a given flavor, $M$, is the unique quantum state formed by superposing the bound state of mirror matter and normal antimatter with the bound state of normal matter and mirror antimatter, such that the particle and the antiparticle with flavor $M$ inside the bound states have the same helicity and chirality, and such that the resulting quantum state for the axion, and hence the axion field, $\phi_{A}$, has the same discrete space-time symmetry properties as the dot product, ${\bf E}\cdot{\bf B}$, ensuring that the Lagrangian density, $\mathcal{L}_{A\gamma}$, is invariant under these transformations (charge-conjugation, space-inversion, and time-reversal; See L. M. Lederman & D. N. Schramm (1989), pg. 121–123). Finally, we combine Eq. (E1) and Eq. (E2) to find the value for the axion-photon coupling strength given in Eq. (7).